OpenStudy (anonymous):

Change 3x + 2y = 10 to slope-intercept form and state the slope and y-intercept. You must show all work.

OpenStudy (anonymous):

slope intercept form is y=mx+c so rearrange x and y on different sides of equality

OpenStudy (anonymous):

3x + 2y = 10 2y = 10 - 3x 2y = -3x +10 y=-(3/2) x +5 so slope is -3/2 and y-intercept is 5 unit

OpenStudy (anonymous):

thanks